# Nonlinear optimals and their role in sustaining turbulence in channel flow

### v1

@@ 4.2 Exploring time-energy parameter space
Figure 10: $T$-$e_{0}$-parameter space. Crosses indicate the quasilinear regime, squares the nonlinear non-localised regime and circles the nonlinear localised regime. The dashed lines were added manually to aid visual distinction between the regimes. Furthermore, each dot or cross is also colour-coded according to its $e_{0}$, which is redundant in this plot but allows establishing connections to Figs. 14, 15(a) and 15(b).
It is clear that not any arbitrary parameter combination of $T$ and $e_{0}$ will give rise to optimals that distil the relevant dynamics of fully-developed turbulence. Although rough estimates for the time and energy scales can be made based on scaling arguments as discussed in Section 4.1, it is important to understand how the shape and dynamics of the optimals change in $(T$,$e_{0})$-space. In particular, we would like to investigate how sensitive the dynamics are to changes in $T$ and $e_{0}$. Figure 10 shows the different regimes that can be identified between $T=0.35h/u_{\tau}$ and $T=2.8h/u_{\tau}$, with the initial energy chosen small enough to result in an end state that is not too chaotic for the optimisation to converge. The gain achieved by each optimal is also shown, with values between the points obtained by interpolation. For small $e_{0}$, the quasilinear regime is found, though the boundary to the nonlinear regime has not been precisely resolved for all $T$ (unlike in the transition problem, this area is not of primary interest here). It is interesting to note, however, that at $T\leq t_{p}$, the quasilinear optimal found at low $e_{0}$ produces higher gains than optimals found at higher $e_{0}$. This trend is reversed for larger $T$, which implies that the primary growth timescale $t_{p}$ can also be thought of as the largest time horizon for which the (quasi-) linear optimal outperforms the nonlinear one.
@@
At large time horizons but small initial energies, a completely new nonlinear optimal structure emerges (Fig. 11). Interestingly, this structure is not localised but global, and, as can be seen in Fig. 12, much of the energy growth happens in the final part of the calculation, which explains why this structure is not optimal when considering shorter time scales. However, as is apparent visually and will be discussed in more detail later, the dynamics of how these optimals evolve bear no resemblance to fully-developed turbulence, which confirms the expectation discussed at the beginning of Section 4 that in a real flow, optimals have to evolve on much shorter time scales. Thus, we can conclude that time scales between about $T=0.7h/u_{\tau}$ and $T=2.1h/u_{\tau}$, and initial energy roughly between $e_{0}/E_{0}={3}\cdot 10^{-5}$ and $e_{0}/E_{0}=10^{-4}$ give rise to nonlinear optimals whose dynamics share distinctive features with fully-developed turbulence.
Figure 13: Early part of the streamwise velocity evolution ($tu_{\tau}/h=0,0.05,0.1$ for comparison with Fig. 4) of the nonlinear optimal ($e_{0}/E_{0}={7}\cdot 10^{-5}$) for $Tu_{\tau}/h=2.1$ plotted in $y$-$z$-plane.
@@
These results are reassuring and at the same time highlight the usefulness of considering nonlinear optimals: whereas linear optimals require a very specific time horizon $T$ in order for their predicted steak spacing to agree with fully-developed turbulence, nonlinear optimals are much more robust in this regard, as every optimal in the nonlinear localised regime, which spans a considerable range of $T$ and $e_{0}$, yields dynamics consistent with fully-developed turbulence.
@@ 5 Conclusion
These results confirm that the evolution of nonlinear optimals reflects crucial aspects of turbulent dynamics. At the same time, the optimisation serves to remove any non-essential dynamics, which would be difficult to filter out in other methods such as DNS. This makes nonlinear optimals an attractive concept for isolating the important dynamics, allowing one to study turbulence in a “clean” environment. At the same time, the present analysis provides upper bounds for the level of energy transfer to the turbulent fluctuations, which could serve as a foundation for future work aimed at better understanding under which circumstances turbulence is sustained. When compared to computationally less expensive linear methods, a major advantage of this approach is the conceptual simplicity and the robustness with which the correct dynamics are found with respect to parameter variations.
@@
Acknowledgements
The work leading to this publication was supported by the PRIME programme of the German Academic Exchange Service (DAAD) with funds from the German Ministry of Education and Research (BMBF).
We thank Sergio Hoyas and Patrick Keuchel for fruitful discussions about the solver implementation.
@@ A.3 Gradient-based optimisation algorithm
We now turn to a more detailed description of the optimisation approach roughly introduced in Section 3. Generally speaking, any gradient-based optimisation algorithm has two components: First, selecting a step direction, and, second, selecting a step size determining how far to move along this direction. Note that in the literature, optimisation problems are often phrased as minimisation problems. Here, since we are trying to maximise the energy gain, we use language to reflect that we are solving a maximisation problem. The mathematics are obviously completely equivalent up to flipping signs. Notable, we refer to gradient ascent rather than gradient descent. A natural choice for the ascent direction is the steepest direction (given by the gradient), however, in the present work, we follow Cherubini et al. (2011) and use the conjugate gradient algorithm with the Polak-Ribière formula (Polak, 1969). In concrete terms, the ascent direction $\mathbf{p}^{n+1}$ is determined using
@@
where $f$ is the objective function and $t$ is a parameter between 00 and $1$, the new guess is either accepted directly, or $\alpha_{i}$ is increased until 18 is no longer satisfied. Note that trying to increase the step size requires additional function evaluations, so in practice, it should usually not be done in every iteration, but only, say, every third iteration. However, not trying to increase the step size once in a while will likely lead to slow convergence due to a possibly too small step size. If 18 is not satisfied, $\alpha_{i}$ is decreased until it is. The step size for the new iteration $\alpha_{i+1}$ is then chosen to be the final step size of the old iteration. Note that upper and lower limits for the step size $\alpha_{i}$ can be imposed to ensure reasonable behaviour of the solver. This was done in the present work, though the borders were chosen quite liberally as a restriction of the steps size should in principle not be required. The convergence criterion is that the relative change of the objective function between two successive iterations lies below a critical value ($10^{-8}$ is chosen in the present work).

### v2

@@ 4.2 Exploring time-energy parameter space
Figure 12: $T$-$e_{0}$-parameter space. Crosses indicate the quasilinear regime, squares the nonlinear non-localised regime and circles the nonlinear localised regime. The dashed lines were added manually to aid visual distinction between the regimes. Furthermore, each dot or cross is also colour-coded according to its $e_{0}$, which is redundant in this plot but allows establishing connections to Figs. 16, 17(a) and 17(b).
It is clear that not any arbitrary parameter combination of $T$ and $e_{0}$ will give rise to optimals that distil the relevant dynamics of fully-developed turbulence. Although rough estimates for the time and energy scales can be made based on scaling arguments as discussed in Section 4.1, it is important to understand how the shape and dynamics of the optimals change in $(T$,$e_{0})$-space. In particular, we would like to investigate how sensitive the dynamics are to changes in $T$ and $e_{0}$. Figure 12 shows the different regimes that can be identified between $T=0.35h/u_{\tau}$ and $T=2.8h/u_{\tau}$, with the initial energy chosen small enough to result in an end state that is not too chaotic for the optimisation to converge. The gain achieved by each optimal is also shown, with values between the points obtained by interpolation. For small $e_{0}$, the quasilinear regime is found, which we here identify by the most energetic wave number mode of the initial optimal disturbance containing most (more than $90\%$, though the exact value does not matter much as long as it is high) of the total energy. The boundary to the nonlinear regime has not been precisely resolved for all $T$ (unlike in the transition problem, this area is not of primary interest here). It is interesting to note, however, that at $T\leq t_{p}$, the quasilinear optimal found at low $e_{0}$ produces higher gains than optimals found at higher $e_{0}$. This trend is reversed for larger $T$, which implies that the primary growth timescale $t_{p}$ can also be thought of as the largest time horizon for which the (quasi-) linear optimal outperforms the nonlinear one.
@@
At large time horizons but small initial energies, a completely new nonlinear optimal structure emerges (Fig. 13). Interestingly, this structure is not localised but global, and, as can be seen in Fig. 14, much of the energy growth happens in the final part of the calculation, which explains why this structure is not optimal when considering shorter time scales. We distinguish this regime from the nonlinear localised regime by checking that the lowest quarter of wave number modes of the initial optimal disturbance contain almost all ($>99.999\%$) of its energy. However, as is apparent visually and will be discussed in more detail later, the dynamics of how these optimals evolve bear no resemblance to fully-developed turbulence, which confirms the expectation discussed at the beginning of Section 4 that in a real flow, optimals have to evolve on much shorter time scales. Thus, we can conclude that time scales between about $T=0.7h/u_{\tau}$ and $T=2.1h/u_{\tau}$, and initial energy roughly between $e_{0}/E_{0}={3}\cdot 10^{-5}$ and $e_{0}/E_{0}=10^{-4}$ give rise to nonlinear optimals whose dynamics share distinctive features with fully-developed turbulence.
Figure 15: Early part of the streamwise velocity evolution ($tu_{\tau}/h=0,0.35,0.7$ for comparison with Fig. 5) of the nonlinear optimal ($e_{0}/E_{0}={7.2}\cdot 10^{-5}$) for $Tu_{\tau}/h=2.1$ plotted in $y$-$z$-plane.
@@
Another interesting point is that the general shape of these optimals closely resembles that of optimal perturbations found in the transition problem. In the parameter study by Farano et al. (2016), both the linear regime and the nonlinear localised regime, which they refer to as highly nonlinear regime, seem to have counterparts that can be observed in the transition problem as well, although the time horizons considered in that study are much longer and the initial energies smaller.
These results are reassuring and at the same time highlight the usefulness of considering nonlinear optimals: whereas linear optimals require a very specific time horizon $T$ in order for their predicted steak spacing to agree with fully-developed turbulence, nonlinear optimals are much more robust in this regard, as every optimal in the nonlinear localised regime, which spans a considerable range of $T$ and $e_{0}$, yields dynamics consistent with fully-developed turbulence, despite the relatively simple equation system employed here.
@@
4.3 Short time high energy optimals
Having expanded the parameter space to very long times, we now also address the question of what happens for very short time horizons. Here, we choose $T=0.1u_{\tau}/h$. In order to compensate for this short time horizon, we select a relatively large initial energy of $e_{0}/E_{0}={4}\cdot 10^{-4}$. The evolution of this optimal, which grows by about an order of magnitude in the given time horizon, is shown in Fig. 18.
][t]1.0 Refer to caption
][t]1.0 Refer to caption
Figure 18: Streamwise velocity evolution ($tu_{\tau}/h=0,0.05,0.1$) of the nonlinear optimal ($e_{0}/E_{0}={4}\cdot 10^{-4}$) for $Tu_{\tau}/h=0.1$ plotted in $x$-$z$-plane (top row) and in the $x$-$y$-plane (bottom row).
Interestingly, there is no localisation in $x$ or $z$, and although the disturbance is concentrated in $y$-direction to where the base shear is high, it appears on both walls, again indicating that localisation is not important for these parameters. This is partly expected due to the relatively high initial energy, but the more important point is probably the short time horizon. As can be inferred from the tilting of the disturbances visible in the bottom row of Fig. 18, only the Orr mechanism is important here, as there is no time for streaks to develop and make use of the lift-up mechanism. This is expected and consistent with the time scales observed in the temporal evolution of the nonlinear optimals discussed in Sections 4.1 and 4.2. The fact that there is only a single linear mechanism at play due to the short time horizon thus helps explain why localisation is not observed.
@@ 5 Conclusion
These results confirm that the evolution of nonlinear optimals reflects crucial aspects of turbulent dynamics, despite the simplicity of our model. At the same time, the optimisation serves to remove any non-essential noise not contributing to disturbance growth, which would be difficult to filter out in other methods such as DNS. This makes nonlinear optimals an attractive concept for isolating the important dynamics, allowing one to study turbulence in a “clean” environment. At the same time, the present analysis provides upper bounds for the level of energy transfer to the turbulent fluctuations, which could serve as a foundation for future work aimed at better understanding under which circumstances turbulence is sustained. When compared to computationally less expensive linear methods, two major advantages of this approach are (i) the conceptual simplicity – in particular there is no need to prescribe a streak amplitude as is the case for methods coupling linear processes “manually” (e.g. Markeviciute & Kerswell, 2024; Ciola et al., 2024) – and (ii) the robustness with which correct values for important characteristics of the dynamics, namely streak spacing and amplitude, are found with respect to parameter variations.
@@
Declaration of interests: The authors report no conflict of interest.
Acknowledgements: We thank Sergio Hoyas and Patrick Keuchel for fruitful discussions about the solver implementation.
Funding: The work leading to this publication was supported by the PRIME programme of the German Academic Exchange Service (DAAD) with funds from the German Ministry of Education and Research (BMBF).
Data availability statement: The data that support the findings of this study are openly available in Pangaea at https://doi.pangaea.de/10.1594/PANGAEA.983358.
@@ A.3 Gradient-based optimisation algorithm
We now turn to a more detailed description of the optimisation approach roughly introduced in Section 3. Generally speaking, any gradient-based optimisation algorithm has two components: First, selecting a step direction, and, second, selecting a step size determining how far to move along this direction. Note that in the literature, optimisation problems are often phrased as minimisation problems. Here, since we are trying to maximise the energy gain, we use language to reflect that we are solving a maximisation problem. The mathematics are obviously completely equivalent up to flipping signs. Notable, we refer to gradient ascent rather than gradient descent. A natural choice for the ascent direction is the steepest direction (given by the gradient), however, in the present work, we follow Cherubini et al. (2011) and use the conjugate gradient algorithm with the Polak-Ribière formula (Polak & Ribiere, 1969). In concrete terms, the ascent direction $\mathbf{p}^{n+1}$ is determined using
@@
where $f$ is the objective function and $t$ is a parameter between 00 and $1$, the new guess is either accepted directly, or $\alpha_{i}$ is increased until 19 is no longer satisfied. Note that trying to increase the step size requires additional function evaluations, so in practice, it should usually not be done in every iteration, but only, say, every third iteration. However, not trying to increase the step size once in a while will likely lead to slow convergence due to a possibly too small step size. If 19 is not satisfied, $\alpha_{i}$ is decreased until it is. The step size for the new iteration $\alpha_{i+1}$ is then chosen to be the final step size of the old iteration. Note that upper and lower limits for the step size $\alpha_{i}$ can be imposed to ensure reasonable behaviour of the solver. This was done in the present work, though the borders were chosen quite liberally as a restriction of the steps size should in principle not be required. The convergence criterion is that the relative change of the objective function between two successive iterations lies below a critical value ($10^{-8}$ is chosen in the present work).